QUANTUM MONTH CARLO PREDICTION OF VIBRATIONAL FREQUENCY SHIFTS OF (He)_{N}-CO_{2} CLUSTERS

\underline{HUI LI} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; NICHOLAS BLINOV and PIERRE-NICHOLAS ROY, Department of Chemistry, University of Alberta, Edmonton, Alberta T6G 2G2, Canada.

High resolution infrared spectra of (He)_{N}-CO_{2} clusters with N up to 60 have been studied by McKellar and co-workers, in the region of the ν_{3} fundamental band of CO_{2}.^{a} By fitting to the rotation-vibration transitions for each cluster size, vibrational band origins ν_{0} and rotational constants B and D were obtained as functions of N. Quantum Monte Carlo simulation predictions of B and D for (He)_{N}-CO_{2} clusters have been found to agree well with experiment for N values up to 17.^{b} However, those simulations were based on two-dimensional potential energy surfaces with CO_{2} fixed at its equilibrium geometry. While an adequate approximation for describing rotational constants, that approach cannot predict the ν_{3} vibrational band origin shift of CO_{2} in (He)_{N} clusters, because of neglect of the dependence of the potential energy surface on the Q_{3} asymmetric-stretch motion of CO_{2}. We recently determined a three-dimensional analytical ‘Morse/Long-Range’ potential energy surface for the CO_{2}-He bimer, which explicitly depends on the Q_{3} asymmetric-stretch vibrational motion of CO_{2}, and also incorporates the correct angle-dependent inverse-power long-range behaviour.^{c} We have used this new potential in path-integral Monte Carlo simulations to predict both the effective rotational constant and the shift of the ν_{3} band origin for CO_{2} doped in (He)_{N} clusters with N up to 100. Our results will be compared with experiment for clusters formed from both symmetric and asymmetric isotopologues of CO_{2}.

a J. Tang et al. Phys. Rev. Lett. 92, 145503 (2004); A.R.W. McKellar, J. Chem. Phys. 128, 044308 (2008).

b F. Paesani, Y. Kwon and K.B. Whaley Phys. Rev. Lett. 94, 153401 (2005); N. Blinov and P.-N. Roy J. Low Temp. Phys.

140, 314 (2005); F. Mezzacapo, J. Low Temp. Phys. 140, 241 (2005).

c H. Li and R.J. Le Roy, PCCP (2008, in press).